# LayerPeeler: Autoregressive Peeling for Layer-wise Image Vectorization

### v1

@@ 3.3. LayerPeeler Dataset
Our annotation process systematically identifies and extracts non-occluded paths within each simplified SVG, denoted as $S$. We define that a path $p$ – representing a shape – is considered non-occluded if no other path both overlaps with it and appears above it in the layer hierarchy. All such topmost paths are collected into a set $\{p_{i}\}_{i=1}^{M}$, where $M$ is the total number of non-occluded paths. Both the complete SVG $S$ and the extracted layer group $\{p_{i}\}_{i=1}^{M}$ are rasterized using a renderer $\mathcal{R}(\cdot)$, yielding images $\mathcal{R}(S)$ and ($\mathcal{R}(\{p_{i}\}_{i=1}^{M})$, respectively. To facilitate accurate VLM-assisted annotation, we design a visual prompting scheme (illustrated in Fig. 4). Specifically, we arrange two images side by side: A (left), the SVG $\mathcal{R}(S)$, providing holistic context, and B (right), the extracted non-occluded layers ($\mathcal{R}(\{p_{i}\}_{i=1}^{M})$, to be described.
The two panels are separated by a vertical black line to aid spatial parsing, and a checkerboard background is used to prevent shapes from blending with the background color. We use Gemini-2.0-Flash, selected for its powerful vision-language performance and cost-efficiency, to generate two types of annotations: semantic descriptions for recognizable elements in earlier stages and geometric descriptions for abstract shapes in later stages. Each annotation is prefixed with "remove" to form an editing instruction $\mathcal{P}_{e}$. For every step, we construct a training triplet $(\mathcal{P}_{e},I_{\mathrm{src}},I_{\mathrm{tar}})$, where $I_{\mathrm{src}}=\mathcal{R}(S)$ is the rasterized image before removal, $I_{\mathrm{tar}}=\mathcal{R}(S^{\prime})$ represents the image after removal with $S^{\prime}=S-\{p_{i}\}_{i=1}^{M}$, and $\mathcal{P}_{e}$ is the corresponding edit prompt. This process iteratively repeats, removing and annotating topmost layers until the SVG is fully decomposed. After a final cleaning phase, we obtain approximately $617,000$ high-quality data triplets, providing rich supervision for training models on fine-grained, layer-wise editing.
Figure 4. The visual annotation used for non-occluded layer captioning.
3.4. Inference
Inference begins by sending the source image to VLMs to get the tuple $\{\mathcal{G},\mathcal{P}_{g},\{\mathcal{B}_{i},\mathcal{P}_{i}\}_{i=1}^{N}\}$ and initializing a random noise vector $z_{T}$, which serves as the starting point for the denoising process. The latents of the source image, denoted as $x_{\mathrm{src}}$, are kept noise-free during denoising steps to provide spatial correspondence and preserve high-frequency details. Once the diffusion model completes the layer-peeling operation, it outputs an image in which the specified topmost layer has been removed. To isolate the removed region, we compute a pixel-level difference map between the source and output images, apply thresholding to create a binary mask, and use this mask to extract the removed region from the source image. The extracted region is then vectorized using an off-the-shelf vectorization tool. This entire pipeline is executed autoregressively: after each layer is removed and vectorized, the resulting output image becomes the new source image for the next iteration. The process continues until the image becomes completely white. Finally, we stack all the vectorized layers in reverse order to reconstruct the complete SVG.
@@ 4. Experiments
4.1. Implementation Details
For training, we initialize the FLUX weights from OmniEditor (Huang et al., 2025), which was pretrained on the SeedEdit (Ge et al., 2024) dataset to provide basic editing capabilities for natural images. We adopt LoRA with a rank of $256$ and a learning rate of $0.0001$. We use learning rate warmup followed by cosine decay and train with a minibatch size of $32$. Training runs for $50,000$ steps on four NVIDIA A100 GPUs using our custom LayerPeeler dataset. During inference, we employ Gemini-2.5-Flash for constructing layer graphs. For image generation, we use $40$ diffusion steps with a guidance scale of $4.5$. For vectorization, we use the tool provided by Recraft***https://www.recraft.ai/, selected for its high-quality output and seamless code integration. However, we emphasize that our approach is agnostic to the specific vectorization method used. In our experiments, we use a test set comprising $250$ randomly selected images rendered from SVG, ensuring no overlap with the training dataset. The average time to vectorize a given image is $128.05$ seconds.
@@ 4.2. Baselines
We compare LayerPeeler against six state-of-the-art rule-based and data-driven methods for layer-wise image vectorization: IVD (Law and Kang, 2025), LIVE (Ma et al., 2022), O&R (Hirschorn et al., 2024), SGLIVE (Zhou et al., 2024), LIVSS (Wang et al., 2024a), and LayerTracer (Song et al., 2025). We exclude several other training-based methods from our comparison due to practical limitations: some methods (e.g., Im2Vec (Reddy et al., 2021)) only handle simple icons and fail on more complex inputs, while others lack a publicly available implementation (Shen and Chen, 2021; Chen et al., 2023; Thamizharasan et al., 2024).
@@ 4.3. Evaluation Metrics
  • Path Semantics. To evaluate the semantic importance of individual shapes, we first obtain captions for each test image using a VLM. We then randomly remove $30\%$ of paths from the SVG and measure the drop in CLIP text-image similarity scores (Radford et al., 2021) between the original and modified renderings. A larger score drop indicates that individual shapes contribute more significantly to the overall semantic meaning.
@@
  • Visual Fidelity. We evaluate how closely the generated SVG matches the source image by computing both Mean Squared Error (MSE) and LPIPS (Zhang et al., 2018) (VGG backbone (Simonyan and Zisserman, 2014)) between the rendered vector graphics and the original image.
@@
[…]
IVD (Law and Kang, 2025) $0.0183$ $69.17$ $0.0135$ $0.0610$
LIVE (Ma et al., 2022) $0.0158$ $96.41$ $0.0033$ $0.0388$
O&R (Hirschorn et al., 2024) $0.0196$ $149.9$ $0.0095$ $0.0694$
SGLIVE (Zhou et al., 2024) $0.0172$ $99.19$ $0.0034$ $0.0415$
LIVSS (Wang et al., 2024a) $0.0075$ $80.58$ $\mathbf{0.0006}$ $0.0099$
LayerTracer (Song et al., 2025) $0.0121$ $74.23$ $0.0508$ $0.0822$
@@ 4.4.1. Quantitative Comparison
Table 1 presents a comprehensive evaluation of all baseline methods. Our LayerPeeler significantly outperforms existing approaches in path semantics and path regularity, reflecting its ability to preserve meaningful shape information and reconstruct smooth, coherent paths. These improvements are primarily attributed to our autoregressive peeling paradigm, which incrementally recovers complete shapes while preserving the logical order of visual layers. In terms of visual fidelity, our method achieves competitive performance relative to the training-free optimization method, LIVSS, that directly minimize MSE. This is made possible by our use of shared positional encodings and localized attention control, which help the model maintain fine-grained alignment with the original image. Overall, the results demonstrate that LayerPeeler achieves strong semantic understanding and visual quality, producing vector graphics that are both structurally coherent and visually faithful to the source image.
@@ 4.4.2. Qualitative Comparison
Fig. 7 presents a side-by-side visual comparison of the outputs from various baseline methods, our LayerPeeler, and the ground truth SVG. To highlight structural differences, we render a black outline along the contour of each shape. (1) The rule-based method IVD lacks semantic awareness during its curvature-guided inpainting process, often resulting in unnatural reconstructions – for example, the distorted angel wing (second column in Fig. 7). (2) For data-driven methods in the training-free category (specifically, LIVE, O&R, SGLIVE), a common strategy is to initialize shapes at the centers of mass of poorly reconstructed regions. However, these methods often lead to excessive use of shapes to approximate the source image, fragmenting semantically meaningful elements and distorting their forms, as evident in all four examples. LIVSS relies on the feature-averaging effect of SDS loss (Poole et al., 2022) to remove high-frequency details and derive abstracted versions of images. It then uses segmentation models to extract the initial SVG. However, this abstraction strategy is less effective on flat-color, highly abstract SVG, resulting in incomplete (e.g., the cake top) or disconnected shapes (e.g., handle of the knife) that deviate from human perception. (3) The training-based method LayerTracer generates a $3\times 3$ image grid to depict a progressive layer construction sequence. However, it does not explicitly model occlusion relationships, which can lead to inconsistent layering – multiple occluded shapes may appear between successive frames. Furthermore, the lack of pixel alignment across frames introduces jagged or uneven boundaries.
@@
Figure 5. Subjective user study results. Our method achieves the highest number of user selections across both visual fidelity and structural coherence.
4.4.3. Subjective User Study
@@
[…]
@@
Table 2. Qualitative results of the ablation study.
Figure 6. Qualitative results of the ablation study. We show the editing prompts above each image and the editing results.
@@ 4.5. Ablation Study
We perform ablation experiments to evaluate the effectiveness of key components in our pipeline. We first examine the importance of the layer graph by removing this component and instead prompting VLMs to directly analyze (check the supplemental material) the set of non-occluded layers. Without the guidance of a layer graph, VLMs often misidentify partially occluded layers as top layers (see Fig. 6), resulting in incorrect layer ordering. Since our diffusion model is trained specifically to operate in non-occluded regions, these misidentifications lead to a degraded output quality, evident in the drop in path regularity and visual fidelity scores in Table 2. Next, we assess the impact of attention-guided control. Disabling this component results in two recurring failure modes: (1) incomplete removal of the intended layer and (2) unintended modifications to unrelated regions (e.g., the purple flower in Fig. 6). Both issues lead to performance degradation across all evaluation metrics. Notably, path semantics remain relatively consistent across all ablations. We attribute this inherent robustness to the design of our progressive peeling pipeline, which enforces a clear and interpretable correspondence between shapes and prompts throughout the process.
@@ 4.6. Evaluation of Generalizability
We provide examples of the autoregressive peeling process for images with different layer complexities in Fig. 8, where the images are sampled from the test set. Additionally, our system demonstrates robust performance and generalizability across a diverse range of SVG styles in Fig. 9. The source images of different styles are downloaded from the Icons8 and Flaticon websites, or are from the test set, and thus are not in the training set.
@@ 5. Conclusion and Discussion
Since our collected dataset only contains flat-color SVG, our system’s performance may degrade when handling semi-transparent elements (see Fig. 11). Additionally, the autoregressive pipeline occasionally fails due to error accumulation when the VLMs and image diffusion model face challenges in accurately following instructions. In the future, we aim to apply data augmentation techniques, such as adding gradient colors and semi-transparent layers, to improve the model’s generalizability. To mitigate error accumulation, we propose implementing an automatic success check after each round of peeling for validation.
@@ Appendix A Overview
This appendix contains supplementary experimental results and implementation details:
  • Additional vectorization examples and qualitative comparisons (Section B);
  • Detailed VLM prompts used throughout our pipeline (Section C).
Appendix B More Results

### v2

@@ 3.4. LayerPeeler Dataset
Our annotation process systematically identifies and extracts non-occluded paths within each simplified SVG, denoted as $S$. We define that a path $p$ – representing a shape – is considered non-occluded if no other path both overlaps with it and appears above it in the layer hierarchy. All such topmost paths are collected into a set $\{p_{i}\}_{i=1}^{M}$, where $M$ is the total number of non-occluded paths. Both the complete SVG $S$ and the extracted layer group $\{p_{i}\}_{i=1}^{M}$ are rasterized using a renderer $\mathcal{R}(\cdot)$, yielding images $\mathcal{R}(S)$ and $\mathcal{R}(\{p_{i}\}_{i=1}^{M})$, respectively. To facilitate accurate VLM-assisted annotation, we design a visual prompting scheme (illustrated in Fig. 4). Specifically, we arrange two images side by side: A (left), the SVG $\mathcal{R}(S)$, providing holistic context, and B (right), the extracted non-occluded layers $\mathcal{R}(\{p_{i}\}_{i=1}^{M})$, to be described.
The two panels are separated by a vertical black line to aid spatial parsing, and a checkerboard background is used to prevent shapes from blending with the background color. We use Gemini-2.0-Flash, selected for its powerful vision-language performance and cost-efficiency, to generate two types of annotations: semantic descriptions for recognizable elements in earlier stages and geometric descriptions for abstract shapes in later stages. Each annotation is prefixed with “remove” to form an editing instruction $\mathcal{P}_{e}$. For every step, we construct a training triplet $(\mathcal{P}_{e},I_{\mathrm{src}},I_{\mathrm{tar}})$, where $I_{\mathrm{src}}=\mathcal{R}(S)$ is the rasterized image before removal, $I_{\mathrm{tar}}=\mathcal{R}(S^{\prime})$ represents the image after removal with $S^{\prime}=S-\{p_{i}\}_{i=1}^{M}$, and $\mathcal{P}_{e}$ is the corresponding edit prompt. This process iteratively repeats, removing and annotating topmost layers until the SVG is fully decomposed. After a final cleaning phase, we obtain approximately $617,000$ high-quality data triplets, providing rich supervision for training models on fine-grained, layer-wise editing.
@@ 4.1. Baselines
We compare LayerPeeler against six state-of-the-art optimization-based and learning-based methods for layer-wise image vectorization: IVD (law2025image), LIVE (ma2022towards), O&R (hirschorn2024optimize), SGLIVE (zhou2024segmentation), LIVSS (wang2024layered), and LayerTracer (song2025layertracer). We exclude several other methods from our comparison due to practical limitations: some methods (e.g., Im2Vec (reddy2021im2vec)) only handle simple icons and fail on more complex inputs, while others lack a publicly available implementation (favreau2017photo2clipart; shen2021clipgen; chen2023editable; thamizharasan2024vecfusion). Please refer to the supplementary material for implementation details and more comparisons (with du2023image).
@@ 4.2. Evaluation Metrics
  • Path Semantics. To evaluate the semantic importance of individual shapes, we first obtain captions for each test image using a VLM. We then randomly remove $30\%$ of paths from the SVG (zhang2024text) and measure the drop in CLIP text-image similarity scores (radford2021learning) between the original and modified renderings. A larger score drop indicates that individual shapes contribute more significantly to the overall semantic meaning.
@@
  • Visual Fidelity. We evaluate how closely the generated SVG matches the source image by computing both Mean Squared Error (MSE) and LPIPS (zhang2018unreasonable) (VGG backbone (simonyan2014very)) between the rendered vector graphics and the original image.
@@
$\big\downarrow$
MSE $\big\downarrow$ LPIPS $\big\downarrow$
IVD (law2025image) $0.0183$ $69.17$ $0.0135$ $0.0610$
LIVE (ma2022towards) $0.0158$ $96.41$ $0.0033$ $0.0388$
O&R (hirschorn2024optimize) $0.0196$ $149.9$ $0.0095$ $0.0694$
SGLIVE (zhou2024segmentation) $0.0172$ $99.19$ $0.0034$ $0.0415$
LIVSS (wang2024layered) $0.0075$ $80.58$ $\mathbf{0.0006}$ $0.0099$
LayerTracer (song2025layertracer) $0.0121$ $74.23$ $0.0508$ $0.0822$
@@ 4.3.1. Quantitative Comparison
Table 1 presents a comprehensive evaluation of all baseline methods. Our LayerPeeler significantly outperforms existing approaches in path semantics and path regularity, reflecting its ability to preserve meaningful shape information and reconstruct smooth, coherent paths. These improvements are primarily attributed to our autoregressive peeling paradigm, which incrementally recovers complete shapes while preserving the logical order of visual layers. In terms of visual fidelity, our method achieves competitive performance relative to the optimization-based method, LIVSS, that directly minimize MSE. This is made possible by our use of shared positional encodings and localized attention control, which help the model maintain fine-grained alignment with the original image. Overall, the results demonstrate that LayerPeeler achieves strong semantic understanding and visual quality, producing vector graphics that are both structurally coherent and visually faithful to the source image.
@@ 4.3.2. Qualitative Comparison
Fig. 8 presents a side-by-side visual comparison of the outputs from various baseline methods, our LayerPeeler, and the ground truth SVG. To highlight structural differences, we render a black outline along the contour of each shape. (1) The optimization-based geometric method, IVD, lacks semantic awareness in its curvature-guided inpainting process, often leading to unnatural reconstructions (e.g., the distorted and incomplete angel wings in the first column of Fig. 8). (2) For optimization-based methods in the differentiable rendering category (LIVE, O&R, SGLIVE), a common strategy is to initialize shapes at the centers of mass of poorly reconstructed regions. However, these methods often lead to excessive use of shapes to approximate the source image, fragmenting semantically meaningful elements, as evident in all four examples. LIVSS relies on the feature-averaging effect of SDS loss (poole2022dreamfusion) to remove high-frequency details and derive abstracted versions of images. It then uses segmentation models to extract the initial SVG. However, this abstraction strategy is less effective on flat-color, highly abstract SVG, resulting in incomplete (e.g., the cake top) or disconnected shapes (e.g., handle of the knife) that deviate from human perception. (3) The learning-based method LayerTracer generates a $3\times 3$ image grid to depict a progressive layer construction sequence. However, it does not explicitly model occlusion relationships, which can lead to inconsistent layering – multiple occluded shapes may appear between successive frames. Furthermore, the lack of pixel alignment across frames introduces jagged boundaries.
@@
Figure 5. Perceptual study results. Our method achieves the highest number of user selections across both visual fidelity and structural coherence.
4.3.3. Perceptual Study
@@
$\big\downarrow$
MSE $\big\downarrow$ LPIPS $\big\downarrow$
Claude-Sonnet-4.0 $0.0232$ $26.65$ $0.0014$ $0.0096$
Qwen2.5-VL-32B-Instruct $0.0219$ $30.56$ $0.0023$ $0.0096$
@@
Figure 6. Qualitative results of the ablation study. We show the editing prompts above each image and the editing results. The icons are from ©SVGRepo.
@@ 4.4. Ablation Study
We perform ablation experiments to evaluate the effectiveness of key components in our pipeline. We first replace the default VLM, Gemini-Flash-2.5, with two alternatives: Claude-Sonnet-4.0 and Qwen-2.5-VL-32B-Instruct, and report the quantitative results in Table 2. The results show that our method remains robust across different VLM choices. Moreover, since Qwen is an open-source model, it provides a practical alternative to proprietary VLMs. We then examine the importance of the layer graph by removing this component and instead prompting VLMs to directly analyze (check the supplemental material) the set of non-occluded layers. Without the guidance of a layer graph, VLMs often misidentify partially occluded layers as top layers (see Fig. 6), resulting in incorrect layer ordering. Since our diffusion model is trained specifically to operate in non-occluded regions, these misidentifications lead to a degraded output quality, evident in the drop in path regularity and visual fidelity scores in Table 2. Next, we assess the impact of attention-guided control. Disabling this component results in two recurring failure modes: (1) incomplete removal of the intended layer and (2) unintended modifications to unrelated regions (e.g., the purple flower in Fig. 6). Both issues lead to performance degradation across all evaluation metrics. Notably, path semantics remain relatively consistent across all ablations. We attribute this inherent robustness to the design of our progressive peeling pipeline, which enforces a clear and interpretable correspondence between shapes and prompts throughout the process.
@@ 4.5. Evaluation of Generalizability
We provide examples of the autoregressive peeling process for images with different layer complexities in Fig. 9. Additionally, our system demonstrates robust performance and generalizability across a diverse range of SVG styles in Fig. 10. The source images in this subsection are downloaded from the Flaticon website, or are from the test set, and thus are not in the training set.
Figure 7. Limitations of our method include the fragmentation of semi-transparent layers into multiple discrete shapes (left panel) and error accumulation during the autoregressive process (right panel), which can produce redundant paths in the final SVG output. The icons are from ©SVGRepo.
@@ 5. Conclusion and Discussion
Since our collected dataset only contains flat-color SVG, our system’s performance may degrade when handling semi-transparent elements (see Fig. 12). Additionally, there are stochastic errors (see supplemental materials for details) in the system arising from incorrect top-layer detection by VLMs and the diffusion model’s inability to accurately remove specified regions, which leads to error accumulation in the autoregressive pipeline. Our method may also struggle with realistic images, primarily because no ground-truth layer-peeling dataset exists for photographs, and VLMs often have difficulty accurately identifying layers in natural images. In the future, we aim to apply data augmentation techniques, such as adding gradient colors, to improve the model’s generalizability. To mitigate error accumulation, we propose implementing an automatic success check after each round of peeling for validation.
Acknowledgements.
The work described in this paper was fully supported by a GRF grant from the Research Grants Council (RGC) of the Hong Kong Special Administrative Region, China [Project No. CityU 11216122].
Figure 8. Qualitative comparison between baselines and LayerPeeler. We render black strokes along the contours of recovered vector elements to better illustrate their structural quality. The rule-based method IVD lacks semantic understanding of each shape; thus, its curvature-based inpainting leads to unnatural completions like the wing of the angel. Training-free methods (particularly LIVE, O&R, and SGLIVE) tend to produce fragmented shapes with irregular components due to their naive depth estimation, as evident in the flower petals. LIVSS’s blurring strategy does not work for non-photorealistic icon images, yielding disjoint or incomplete shapes (e.g., the head of the angel). LayerTracer trains a model to predict the SVG construction sequence, but the alignment between two frames is not guaranteed. In contrast, LayerPeeler generates paths that are both semantically meaningful and geometrically regular.
Figure 9. Additional examples demonstrating our autoregressive peeling process on images with varying layer complexity.
Figure 10. Evaluation of generalizability across different styles of source images, including flat-color icons, pixel art, and stickers.
Figure 11. Additional examples demonstrating the construction and progressive updates of layer graphs throughout the peeling process.
Figure 12. Limitations of our method include the fragmentation of semi-transparent layers into multiple discrete shapes (left panel) and error accumulation during the autoregressive process (right panel), which can produce redundant paths in the final SVG output.
@@ Appendix A Overview
This appendix contains supplementary implementation details and experimental results.
Appendix B Implementation Details
For training, we initialize the FLUX weights from OmniEditor (huang2025photodoodle), which was pretrained on the SeedEdit (ge2024seed) dataset to provide basic editing capabilities for natural images. We adopt LoRA with a rank of $256$ and a learning rate of $0.0001$. We use learning rate warmup followed by cosine decay and train with a minibatch size of $32$. Training runs for $50,000$ steps on four NVIDIA A100 GPUs using our custom LayerPeeler dataset. During inference, we employ Gemini-2.5-Flash for constructing layer graphs. For image generation, we use $40$ diffusion steps with a guidance scale of $4.5$. For vectorization, we use the tool provided by Recraft***https://www.recraft.ai/, selected for its high-quality output and seamless code integration. However, we emphasize that our approach is agnostic to the specific vectorization method used. In our experiments, we use a test set comprising $250$ randomly selected images rendered from SVG, ensuring no overlap with the training dataset.
Appendix C Runtime Analysis
On an NVIDIA RTX $5880$ Ada GPU, the average vectorization time per image is $128.05$ seconds. Each peeling iteration takes approximately $26.2$s for the VLM, $15$s for the diffusion model, and $4.7$s for vectorization. Notably, the number of iterations required for complex images is often comparable to that for simpler ones: images with more paths may have similar (or even lower) layer complexity than those with fewer paths. As a result, the overall runtime remains relatively stable. Our pipeline is thus faster than optimization-based approaches, which typically require over $10$ minutes per image.
Appendix D More Comparisons
du2023image proposed an optimization-based geometric method that takes a user-provided segmentation map as input and generates an SVG with gradient colors and semi-transparent layers. To recover occluded regions, the method assumes that upper layers are semi-transparent. Consequently, it is limited to simple overlap scenarios, depends heavily on per-pixel segmentation masks, and lacks a dedicated shape-completion step. In other words, while it can merge visible regions in the segmentation map to form a complete shape, it cannot reconstruct fully invisible regions, since these never appear in the segmentation maps.
When comparing LayerPeeler with their method, we encountered a known issue in the released codehttps://github.com/Zhengjun-Du/ImageVectorViaLayerDecomposition/issues/1 that prevents the generation of the final SVG output. As this issue remains unresolved, only $195$ out of $250$ input images successfully produced SVG results. Therefore, we omit a direct comparison in the main paper and instead provide a side-by-side comparison in Fig. 13. The results show that while their method benefits from segmentation maps to produce clean shape boundaries, the absence of a true shape completion step leads to incomplete reconstructions (e.g., the top of the cake and the angel’s wing). In practice, their method can only handle simple overlaps (e.g., the blade of the dagger), whereas our method explicitly targets the reconstruction of invisible regions.
We also report quantitative results of du2023image on the $195$ test images in Table 3. The Path Irregularity score indicates that their method fails to generate shapes consistent with the ground-truth SVG due to the limitation of their method.
Figure 13. Qualitative Comparison with du2023image. At first glance, their method appears to produce clean outlines, but this actually reflects incomplete shapes (i.e., no reconstruction occurs in invisible regions that are not indicated by the segmentation map). The icons are from ©SVGRepo.
Table 3. Quantitative comparison with du2023image.
Method
Path
Semantics
$\big\uparrow$
Path
Irregularity
$\big\downarrow$
MSE $\big\downarrow$ LPIPS $\big\downarrow$
du2023image $0.0216$ $44.08$ $0.0043$ $0.0173$
LayerPeeler (Ours) $\mathbf{0.0242}$ $\mathbf{25.41}$ $\mathbf{0.0011}$ $\mathbf{0.0083}$
Appendix E Complex Images
Our primary test set consists mainly of flat-color icons. To provide a fairer evaluation, we additionally collect a set of $20$ complex stickers (averaging $46$ paths, compared to $10.8$ in the original test set) and evaluate all baselines on them. As image complexity increases, LayerPeeler consistently outperforms competing methods (see Table 4). In contrast, optimization-based approaches (e.g., LIVE, O&R, and SGLIVE) can process complex images by introducing a large number of paths, but their outputs often suffer from fragmented shapes and disorganized layer structures, as reflected in the Path Semantics and Path Irregularity metrics. Moreover, Fig. 14 shows that the visual trends observed with simple icons persist in these more complex cases. Overall, these results demonstrate LayerPeeler’s ability to preserve coherent structure and semantic organization even for highly complex inputs. Notably, we exclude du2023image from this comparison, since its runtime grows exponentially with the number of regions, preventing it from producing results within a reasonable time.
Table 4. Quantitative comparison over complex images.
Method
Path
Semantics
$\big\uparrow$
Path
Irregularity
$\big\downarrow$
MSE $\big\downarrow$ LPIPS $\big\downarrow$
IVD (law2025image) $0.0194$ $28.97$ $0.0083$ $0.0319$
LIVE (ma2022towards) $0.0191$ $42.18$ $0.0074$ $0.0633$
O&R (hirschorn2024optimize) $0.0198$ $46.14$ $0.0229$ $0.1214$
SGLIVE (zhou2024segmentation) $0.0187$ $42.33$ $0.0088$ $0.0622$
LIVSS (wang2024layered) $0.0099$ $30.71$ $\mathbf{0.0014}$ $0.0270$
LayerTracer (song2025layertracer) $0.0118$ $36.11$ $0.0426$ $0.1121$
LayerPeeler (Ours) $\mathbf{0.0244}$ $\mathbf{9.733}$ $0.0026$ $\mathbf{0.0149}$
Figure 14. Qualitative comparison of complex images. These results demonstrate that LayerPeeler generalizes well to more complex stickers. The sticker images are from ©Flaticon.
Appendix F Stochastic Errors
Our LayerPeeler is built upon powerful VLMs and image diffusion models, both of which can introduce stochastic errors during execution. To quantify these errors, we sample $30$ icon images from the test set and vectorize each image three times using different random seeds. Across a total of $318$ peeling iterations, we observe 0 invalid JSON errors, $8$ incorrect top-layer detection errors, and $19$ diffusion model errors. These results indicate that the diffusion process is the primary bottleneck of our pipeline. To alleviate stochastic errors, our method already employs several strategies, including layer graph representation and updating, local/global attention control, and a positional encoding scheme. Furthermore, our approach is not restricted to a specific model version: the vectorization paradigm can be adapted to more advanced models (e.g., the recently introduced Flux.1-Kontext), potentially yielding more robust performance and lower error rates.
Appendix G More Results